OpenStudy (johnnydicamillo):

Consider the function f(X) whose second derivative is f''(x) = 2x + 4sin(x). If f(0) = 3 and f'(0) = 4, what is f(x)

OpenStudy (johnnydicamillo):

so I know the anti derivative of the function is \[x^2 -4\cos(x) + c\]

OpenStudy (johnnydicamillo):

Now we need to find that constant knowing that f'(0) = 4 so we set it up like\[f(0) = 0^2 -4\cos(0) + c = 3\]

OpenStudy (johnnydicamillo):

scratch that it = 4 so in this case c = 0

OpenStudy (anonymous):

f'(x) = x ^2- 4cos(x)+c not f(x)

OpenStudy (johnnydicamillo):

okay

OpenStudy (johnnydicamillo):

then I find the derivative of that f(x) is \[\frac{ x^3 }{ 3 } - 4\sin(x) + c\]

OpenStudy (anonymous):

\[f^{'}(0)=4\] \[f^{'}(0)=(0)^2-4cos(0)+c=4\]

OpenStudy (anonymous):

find C first and then integrate

OpenStudy (johnnydicamillo):

c = 0

OpenStudy (johnnydicamillo):

-4 * 1 = 4

OpenStudy (johnnydicamillo):

wait

OpenStudy (johnnydicamillo):

8?

OpenStudy (anonymous):

\[c=8\] \[f^{'}(x)=x^2-4cos(x)+8\]

OpenStudy (anonymous):

correct

OpenStudy (johnnydicamillo):

right right, then we get \[f(x) = \frac{ x^3 }{ 3 } - 4\sin(x) + 8x + c\]

OpenStudy (anonymous):

correct

OpenStudy (johnnydicamillo):

so c = 3 because everything 0s out

OpenStudy (anonymous):

yes

OpenStudy (johnnydicamillo):

thanks for catching that error

OpenStudy (anonymous):

np